OpenStudy (anonymous):

If a can of ravioli with a diameter of 2 inches and a height of 4.5 inches holds 12 ounces of ravioli, how many ounces of ravioli would a can with the same height, but double the diameter, hold?

OpenStudy (campbell_st):

r = 1inch (diameter 2) \[v = \pi \times1^2 \times4.5\] then \[V = 4.5\pi\] now consider r = 2 (diameter 4) \[V = \pi \times 2^2 \times 4.5\] \[V = 18\pi\] the volume is 4 times the original so you would expect 4 x 12 = 48 ounces of ravioli